# Benchmarking Unsupervised Online IDS for Masquerade Attacks in CAN

### v1

@@ Abstract
Vehicular controller area networks (CANs) are susceptible to masquerade attacks by malicious adversaries. In masquerade attacks, adversaries silence a targeted ID and then send malicious frames with forged content at the expected timing of benign frames. As masquerade attacks could seriously harm vehicle functionality and are the stealthiest attacks to detect in CAN, recent work has devoted attention to compare frameworks for detecting masquerade attacks in CAN. However, most existing works report offline evaluations using CAN logs already collected using simulations that do not comply with domain’s real-time constraints. Here we contribute to advance the state of the art by introducing a benchmark study of four different non-deep learning (DL)-based unsupervised online intrusion detection systems (IDS) for masquerade attacks in CAN. Our approach differs from existing benchmarks in that we analyze the effect of controlling streaming data conditions in a sliding window setting. In doing so, we use realistic masquerade attacks being replayed from the ROAD dataset. We show that although benchmarked IDS are not effective at detecting every attack type, the method that relies on detecting changes at the hierarchical structure of clusters of time series produces the best results at the expense of higher computational overhead. We discuss limitations, open challenges, and how the benchmarked methods can be used for practical unsupervised online CAN IDS for masquerade attacks.
@@
Streaming Data, Online Algorithms, Machine Learning, Anomaly Detection, In-Vehicle Security, CAN Bus
1 Introduction
@@
Today’s vehicles are intricate cyber-physical systems with growing connectivity capabilities to interact with vehicle-to-everything (V2X) technologies [1, 2, 3]. This increasing connectivity expands the attack surface to adversaries that can access in-vehicle networks. The Controller Area Network (CAN) is the most prevalent protocol used on in-vehicle networks allowing hundreds of electronic control units (ECUs) communicate vehicle functionality [4, 5, 6]. While the lack of security features makes CAN lightweight and inexpensive, it is susceptible to attacks that aim to manipulate communications and significantly degrade in-vehicle network performance. This has been shown to result in life threatening incidents such as unintended acceleration, brake manipulation, and rogue steering wheel turning [7, 8].
@@
CAN attacks are usually classified as fabrication, suspension, and masquerade attacks [9, 10]. From these, masquerade attacks are the hardest to detect because they do not perturb the regular frame timing which can be detected using time-based methods [11, 12]. Instead, in masquerade attacks, the adversary first silences a targeted ECU and then injects spoofed frames which alters the content of regular frames instead of time-related patterns. This makes masquerade attacks the stealthiest often having an effect on the regular evolution of vehicle states usually measured by sensors [13]. Therefore, recent research efforts have been channeled on designing and deploying intrusion detention systems (IDS) against advanced masquerade attacks in CAN [14, 15, 16, 17].
@@
IDS for detecting masquerade attacks in CAN are often based on unsupervised anomaly detection techniques aiming to characterize the regular relationships between physical signals in a vehicle [15, 18]. Some of them focus on mining time series relationships based on deep learning (DL) models that consequently have high computational cost [19, 16, 20]. In addition, most of existing CAN IDS approaches are tested offline using datasets or CAN logs already collected from a real or simulated environment. This has been synthesized in existing benchmarking frameworks and comparative studies in the CAN IDS space [21, 18]. The offline evaluation approach is different from online evaluation in that the latter is performed using streaming CAN data from a vehicle, simulation, or data log replay [22]. Therefore, offline evaluation hinders the understanding of realistic capabilities and limitations of CAN IDS in a streaming environment. By benchmarking non DL-based unsupervised online IDS for masquerade attacks in CAN, we seek to provide insights intro the effectiveness and limitations of these methods in resource-constrained and streaming environments.
@@
In this paper, we address this gap of knowledge by introducing an empirical and unbiased benchmark of four different non-DL-based unsupervised online IDS for masquerade attacks: Matrix Collelation Distribution, Matrix Correlation Correlation, Ganesan17 [23], and Moriano22 [15]. Here, Matrix Correlation Distribution and Matrix Correlation Correlation methods are simple proposed baselines while Ganesan17 and Moriano22 are already proposed methods in the literature. We test them in simulated masquerade attacks from the ROAD dataset [10]. We rely on a sliding window approach to simulate data log replays of a continuous data stream [24, 25]. We analyze the joint effect of window size and offset on the AUC-ROC as evaluation metric for intrusion detection across the different attacks in the ROAD dataset. We also report an analysis of the testing time per window (TTW) [20] of CAN IDS as a proxy of inference time and discuss insights into the real-time detection capability of these approaches. We further contribute to advance the state of the art in this topic by making publicly available the code of the IDS considered in this study [26]. This furthers research in this topic by allowing other researchers to replicate our results and compare new algorithms with the benchmarked IDS.
@@ 2.1 CAN Protocol
The arbitration ID serves as a unique identifier for messages transmitted over the bus. This ID is crucial for distinguishing between different messages and plays a key role in determining message priority. Lower ID values indicate higher message priority, allowing critical messages to be sent without delay. The data field contains the actual information being transmitted between ECUs. The payload consists of up to 8 bytes of data. Each or multiple bytes represent a specific piece of information, known as “signal”, such as sensor readings or status information. The data payload is essential for exchanging real-time data between ECUs. CAN frames with a certain arbitration ID are transmitted at a predetermined frequency with fixed payload structure. For instance, the power control module transmits ID 0x102 with data field containing engine RPM, vehicle speed, and odometer signals every 0.05s [10].
@@
2.2 CAN Vulnerabilities
@@
In terms of attacks that manipulate CAN messages, there are three categories of exploits:
  • Fabrication attacks, employs a compromised ECU to inject messages with malicious IDs and data fields. Typically CAN IDs are deliberately set to low values, signifying high priority. Continuously injecting such high-priority messages can eventually disrupt legitimate ECUs from transmitting theirs leading to DoS-type attacks or targeting vehicle functions.
  • Suspension attacks, where the adversary seeks to compromise an ECU to prevent it from transmitting some or all of its messages. This could be achieved by removing an ID frame(s) to disconnect the ECU from the CAN bus, thereby halting the transmission of corresponding signal values.
  • Masquerade attacks, the most sophisticated category, involve the adversary initially suspending messages from a specific ID of a targeted ECU, then spoofing messages using this ID at a realistic frequency, thus effectively masquerading as the compromised ECU.
3 Related Work
Here we review and contrast previous offline benchmark studies on CAN IDS (Section 3.1) and online CAN IDS methods (Section 3.2).
3.1 Prior Work Closely Related to the Present Study
@@
3.2 Other Prior Work Related to the Present Study
@@
4 Methods
@@
Note that although streaming and online algorithms are often used as synonyms, here we make the distinction between these two and point out that streaming algorithms process each point as soon as it is received, while online algorithms process small batches at a time [25, 40, 41]. In our work, we use the latter definition, when we refer to online algorithms, i.e., algorithms that repeatedly process small parts of a continuous time series stream. Note that our inclination to benchmark unsupervised learning methods is guided in part by the fact that it is challenging to apply supervised methods on streaming data due to their static nature [42].
4.1 Threat Model
We assume a scenario where an adversary has already gained access to the CAN bus (e.g., by the OBD-II port or through wireless communication) [43]. Here, the adversary compromises two ECUs, one as a strong attacker (i.e., fully compromising an ECU including capabilities for injecting arbitrary attack messages) and the other as a weak attacker (i.e., being able to stop/suspend the ECU from transmitting certain messages) [9].
4.2 Time Windows
@@ 4.3.1 Training Phase (P1):
We fit data-driven models for each of the benchmarked methods using the training data, i.e., $\mathbf{X}_{train}$. Fitted models in this task are then used as the regular (normal) state representation from which we derive comparisons with models fitted at the testing stage. The details of the fitted models for each of the becnhmarked methods are in Section 4.4.
4.3.2 Testing Phase (P2):
The testing phase executes similar tasks as already discussed in the training phase (see tasks T1 and T2 in Section 4.3.1) and adds T3. Note that while P1 is run only once, the tasks in P2 are run continuously and sequentially as new messages are processed and there is a need to check for anomalies. We detail here changes with respect to specific parameters.
@@ 4.3.2 Testing Phase (P2):
In the testing stage, data is processed using a sliding window approach (see more details in Section 4.2). Specifically, smaller sections of $\mathbf{X}_{test}$ of length $\omega$ are selected to create a view of the steaming ($\in\mathbb{R}^{\omega\times n}$). Benchmarked methods consume this batch of data to arrive at conclusions.
@@
We fit the benhmarked methods with their specific models on $\mathbf{X}_{test}$. The specifics of the fitted models are discussed in Section 4.4.
@@
This step entails a data-driven decision-making for deciding if a particular window is containing an anomaly or not. The specifics of the criteria for detection are discussed in Section 4.4.
4.4 Benchmarked Algorithms
This subsection describes in detail each of the four algorithms used in this research. For the sake of readability, we associate each algorithm name of the first author and the last two digits of the publication year (when applicable). Specifically, we describe the following methods: Matrix Correlation Distribution (see Section 4.4.1), Matrix Correlation Correlation (see Section 4.4.2), Ganesan17 [23] (see Section 4.4.3), and Moriano22 [15] (see Section 4.4.4).
@@ 4.4 Benchmarked Algorithms
4.4.1 Implementation of Matrix Correlation Distribution
@@
4.4.2 Implementation of Matrix Correlation Correlation
@@
4.4.3 Implementation of Ganesan17
This method focuses on computing clusters of time series defining a context (such as aggressive or sudden driving). Note that this method does not rely on cluster characterization during training and only relies on processing testing windows. Specifically, for each time window of length $\omega$, the DBSCAN clustering method [52] is applied to a distance matrix $\mathbf{D}_{test}\in\mathbb{R}^{n\times n}$ computed as $2*(1-\mathbf{R}_{test})$, i.e., a squared Euclidean distance to identify clusters [53]. DBSCAN is a density-based non-parametric clustering algorithm that works by grouping points with many nearby neighbors and marking as outliers the points that lie alone in low-density regions. By analyzing clusters of time series as opposed to individual pairs, this method aims to capture richer relationships between signals. We used the same parameter values used in Ganesan et al. [23], namely $eps=1$ capturing how close points should be to each other to be considered part of a cluster and $min\_samples=1$ capturing the minimum number of points to form a dense region.
@@ 4.4.4 Implementation of Moriano22
For detection, in each time window of length $\omega$, we compute the clustering similarity $s$ between the hierarchical clustering from training (computed once) and the hierarchical clustering computed during that specific time window. Here, similarity is used as a proxy of the distance between both hierarchical clusterings. We compute the similarity between hierarchical clustering using the CluSim method [55] based on the work by Gates et al. [56]. We parametrized CluSim by letting $r=-5.0$ and $\alpha=0.9$. Here, $r$ is a scaling parameter that defines the relative importance of memberships at different levels of the hierarchy. That is, the larger $r$, the more emphasis on comparing lower levels of the dendrogram (zoom in). In addition, $\alpha$ is a parameter that controls the influence of hierarchical clusterings with shared lineages. That is, the larger $\alpha$, the further the process will explore from the focus data element, so more of the cluster structure is taken into account into the comparison. The similarity value provided by this method, $s$, lies in the range $[0,1]$, where 0 implies maximally dissimilar clusters, and 1 corresponds to identical clusters. Therefore, we report the anomaly score (or probability of detection) as $1-s$.
4.5 Dataset

### v2

@@ Abstract
Vehicular controller area networks (CANs) are susceptible to masquerade attacks by malicious adversaries. In masquerade attacks, adversaries silence a targeted ID and then send malicious frames with forged content at the expected timing of benign frames. As masquerade attacks could seriously harm vehicle functionality and are the stealthiest attacks to detect in CAN, recent work has devoted attention to compare frameworks for detecting masquerade attacks in CAN. However, most existing works report offline evaluations using CAN logs already collected using simulations that do not comply with the domain’s real-time constraints. Here we contribute to advance the state of the art by introducing a benchmark study of four different non-deep learning (DL)-based unsupervised online intrusion detection systems (IDS) for masquerade attacks in CAN. Our approach differs from existing benchmarks in that we analyze the effect of controlling streaming data conditions in a sliding window setting. In doing so, we use realistic masquerade attacks being replayed from the ROAD dataset. We show that although benchmarked IDS are not effective at detecting every attack type, the method that relies on detecting changes in the hierarchical structure of clusters of time series produces the best results at the expense of higher computational overhead. We discuss limitations, open challenges, and how the benchmarked methods can be used for practical unsupervised online CAN IDS for masquerade attacks.
@@
Streaming Data, Online Algorithms, Machine Learning, Anomaly Detection, In-Vehicle Security, CAN Bus.
I Introduction
@@
Today’s vehicles are intricate cyber-physical systems with growing connectivity capabilities to interact with vehicle-to-everything (V2X) technologies [1, 2, 3]. This increasing connectivity expands the attack surface to adversaries that can access in-vehicle networks. The Controller Area Network (CAN) is the most prevalent protocol used on in-vehicle networks allowing hundreds of electronic control units (ECUs) communicate vehicle functionality [4, 5, 6]. While the lack of security features makes CAN lightweight and inexpensive, it also makes it susceptible to attacks that aim to manipulate communications and significantly degrade in-vehicle network performance. This has been shown to result in life threatening incidents such as unintended acceleration, brake manipulation, and rogue wheel steering [7, 8].
@@
CAN attacks are usually classified as fabrication, suspension, and masquerade attacks [9, 10]. From these, masquerade attacks are the hardest to detect because they do not perturb the regular frame timing which can be detected using time-based methods [11, 12]. Instead, in masquerade attacks, the adversary first silences a targeted ECU and then injects spoofed frames which alters the content of regular frames instead of time-related patterns. This makes masquerade attacks the stealthiest, often having an effect on the regular evolution of vehicle states usually measured by sensors [13]. Therefore, recent research efforts have been channeled on designing and deploying intrusion detection systems (IDS) against advanced masquerade attacks in CAN [14, 15, 16, 17].
@@
IDS for detecting masquerade attacks in CAN are often based on unsupervised anomaly detection techniques aiming to characterize the regular relationships between physical signals in a vehicle [15, 18]. Some of them focus on mining time series relationships based on deep learning (DL) models that consequently have high computational cost [19, 16, 20]. In addition, most existing CAN IDS approaches are tested offline using datasets or CAN logs already collected from a real or simulated environment. This has been synthesized in existing benchmarking frameworks and comparative studies in the CAN IDS space [21, 18]. The offline evaluation approach is different from online evaluation, in that the latter is performed using streaming CAN data from a vehicle, simulation, or data log replay [22]. Therefore, offline evaluation hinders the understanding of realistic capabilities and limitations of CAN IDS in a streaming environment. By benchmarking non DL-based unsupervised online IDS for masquerade attacks in CAN, we seek to provide insights into the effectiveness and limitations of these methods in resource-constrained and streaming environments.
@@
In this paper, we address this gap of knowledge by introducing an empirical and unbiased benchmark of four different non-DL-based unsupervised online IDS for masquerade attacks: Matrix Collelation Distribution, Matrix Correlation Correlation, Ganesan17 [23], and Moriano22 [15]. We focus on non DL-based approaches because vehicular data change over time and space. This dynamic nature makes it difficult for DL-based methods to extract relevant features, often leading to poor performance [24]. Here, Matrix Correlation Distribution and Matrix Correlation Correlation methods are simple proposed baselines while Ganesan17 and Moriano22 are already proposed methods in the literature. We test them in simulated masquerade attacks from the ROAD dataset [10]. We rely on a sliding window approach to simulate data log replays of a continuous data stream [25, 26]. We analyze the joint effect of window size and offset on the AUC-ROC as evaluation metric for intrusion detection across the different attacks in the ROAD dataset. We also report an analysis of the testing time per window (TTW) [20] of CAN IDS as a proxy of inference time and discuss insights into the real-time detection capability of these approaches. We further contribute to advance the state of the art in this topic by making publicly available the code of the IDS considered in this study [27]. This furthers research in this topic by allowing other researchers to replicate our results and compare new algorithms with the benchmarked IDS.
@@ II-A CAN Protocol
CAN is physically implemented as two wires. When a ECU sends a 1 bit, the voltage of the two wires are left unchanged, whereas a 0 bit is sent by manipulating the voltage of one wire up and the other down. Since this is a broadcast network (i.e., all ECUs can talk and listen on the two wires), if one ECU sends a 0, it will overwrite any 1 being sent. Hence, arbitration, a process by which the ECUs coordinate who gets to talk is needed to avoid conflicts. The arbitration ID identifies messages on the CAN bus [9, 10]. When multiple ECUs broadcast simultaneously, they send the ID value of their CAN frame one bit at a time, starting with the most significant bit. ECUs sending a 1 are overwritten by any ECUs sending a 0, and these overwritten ECUs cease their transmission. These ECUs then withdraw from contention and switch to receive mode. The ECU with the lowest ID wins arbitration and transmits its message. ECUs that lose arbitration retry transmission when the bus becomes available.
The data field contains the actual information being transmitted between ECUs. The payload consists of up to 8 bytes of data. Each or multiple bytes represent a specific piece of information, known as “signals,” such as sensor readings or status information. The data payload is essential for exchanging real-time data between ECUs. CAN frames with a certain arbitration ID are transmitted at a predetermined frequency with fixed payload structure. For instance, the power control module transmits ID 0x102 with data field containing engine RPM, vehicle speed, and odometer signals every 0.05s [10].
@@
II-B Threat Model
We assume a scenario where an adversary has already gained access to the CAN bus (e.g., by the OBD-II port or through wireless communication) [29]. Here, the adversary may compromise two ECUs, one as a strong attacker (i.e., fully compromising an ECU including capabilities for injecting arbitrary attack messages) and the other as a weak attacker (i.e., weakly compromising an ECU including capabilities to stop/suspend the ECU from transmitting certain messages) [9]. Although it is easier for an adversary to become a weak rather than a stronger attacker, researchers have demonstrated how to become strong attackers [30, 31].
II-C CAN Vulnerabilities
@@
In terms of attacks that manipulate CAN messages, there are three main types [9, 10]:
  • Fabrication attacks: A strong attacker uses a compromised ECU to inject messages with malicious IDs and data fields. These IDs are often set to low values, which indicate high priority. Repeatedly injecting high-priority messages can block legitimate ECUs, causing denial-of-service (DoS) attacks or disrupting vehicle functions.
  • Suspension attacks: A weak attacker compromises an ECU to stop it from sending some or all of its messages. This is done by removing ID frames, disconnecting the ECU from the CAN bus. This not only affects the compromised ECU but also disrupts receiver ECUs relying on those messages.
  • Masquerade attacks: This advanced attack requires two compromised ECUs, one as a strong attacker and one as a weak attacker. The weak attacker monitors the frequency of a targeted ID and stops the legitimate ECU once it learns the frequency. The strong attacker then injects messages with the same ID at the original frequency. The ID frequency remains unchanged, but the source changes, and the messages contain forged content.
We distinguish between spoofing and masquerade attacks. In spoofing attacks, a compromised ECU sends false (spoofed) frames to trigger a reaction from legitimate ECUs, but the legitimate ECU continues to function. Existing datasets, such as HCRL Car Hacking [33] and can-train-and-test [34], include spoofing/masquerade attacks that overwhelm the CAN bus with both spoofed and legitimate messages, referred to as unsophisticated masquerade attacks [34]. These attacks make detection easier because the frequency of the targeted ID changes, but legitimate signals are not suppressed. As a result, simpler methods based on frequency detection can be used [35]. Our work focuses on sophisticated masquerade attacks from the ROAD dataset [10], the only dataset, to our knowledge, that releases such attacks where legitimate signals are suppressed.
III Related Work
Here we review and contrast previous offline benchmark studies on CAN IDS (Section III-A) and online CAN IDS methods (Section III-B).
III-A Prior Work Closely Related to the Present Study
@@
III-B Other Prior Work Related to the Present Study
@@
Marfo et al. [46] introduced a graph ML framework for detecting masquerade attacks in the CAN bus. They combined shallow graph embeddings with time series features to enhance the detection of masquerade attacks. They show that this approach leads to statistically significant improvements in the detection rates of masquerade attacks compared to a baseline that uses only graph-based features.
IV Methods
@@
Although the terms streaming and online algorithms are sometimes used interchangeably in the literature, we draw a distinction in line with prior surveys on data stream mining and adaptive learning [26, 47, 48]. In our context, streaming algorithms refer to those that process each data instance individually as it arrives, typically under strict memory and time constraints. In contrast, we use online algorithms to describe methods that process small batches or windows of data incrementally, allowing for limited buffering and adaptation over time. Throughout this work, we adopt the latter interpretation. When we refer to online algorithms, we mean algorithms designed to operate on continuously arriving data by processing small, sequential segments of the stream, updating models and decisions incrementally. Note that our inclination to benchmark unsupervised learning methods is guided in part by the fact that it is challenging to apply supervised methods on streaming data due to their static nature [49].
IV-A Time Windows
@@ IV-B1 Training Phase (P1)
We fit data-driven models for each of the benchmarked methods using the training data, i.e., $\mathbf{X}_{train}$. Fitted models in this task are then used as the regular (normal) state representation from which we derive comparisons with models fitted at the testing stage. The details of the fitted models for each of the becnhmarked methods are in Section IV-C.
IV-B2 Testing Phase (P2)
The testing phase executes similar tasks as already discussed in the training phase (see tasks T1 and T2 in Section IV-B1) and adds T3. Note that while P1 is run only once, the tasks in P2 are run continuously and sequentially as new messages are processed and there is a need to check for anomalies. We detail here changes with respect to specific parameters.
@@ IV-B2 Testing Phase (P2)
In the testing stage, data is processed using a sliding window approach (see more details in Section IV-A). Specifically, smaller sections of $\mathbf{X}_{test}$ of length $\omega$ are selected to create a view of the steaming ($\in\mathbb{R}^{\omega\times n}$). Benchmarked methods consume this batch of data to arrive at conclusions.
@@
We fit the benhmarked methods with their specific models on $\mathbf{X}_{test}$. The specifics of the fitted models are discussed in Section IV-C.
@@
This step entails a data-driven decision-making for deciding if a particular window is containing an anomaly or not. The specifics of the criteria for detection are discussed in Section IV-C.
IV-C Benchmarked Algorithms
This subsection describes in detail each of the four algorithms used in this research. Note that we focus on non DL-based algorithms because the dynamic nature of CAN data makes it hard for DL-based methods to extract useful features, which reduces their performance [24]. For the sake of readability, we associate each algorithm name of the first author and the last two digits of the publication year (when applicable). Specifically, we describe the following methods: Matrix Correlation Distribution (see Section IV-C1), Matrix Correlation Correlation (see Section IV-C2), Ganesan17 [23] (see Section IV-C3), and Moriano22 [15] (see Section IV-C4).
@@ IV-C Benchmarked Algorithms
Note that the benchmarked algorithms suit streaming environments. They process small data batches and do not rely on large, representative training datasets, unlike DL-based methods.
IV-C1 Implementation of Matrix Correlation Distribution
@@
IV-C2 Implementation of Matrix Correlation Correlation
@@
IV-C3 Implementation of Ganesan17
This method focuses on computing clusters of time series defining a context (such as aggressive or sudden driving). Note that this method does not rely on cluster characterization during training and only relies on processing testing windows. Specifically, for each time window of length $\omega$, the DBSCAN clustering method [58] is applied to a distance matrix $\mathbf{D}_{test}\in\mathbb{R}^{n\times n}$ computed as $2*(1-\mathbf{R}_{test})$, i.e., a squared Euclidean distance to identify clusters [59]. DBSCAN is a density-based non-parametric clustering algorithm that works by grouping points with many nearby neighbors and marking as outliers the points that lie alone in low-density regions. By analyzing clusters of time series as opposed to individual pairs, this method aims to capture richer relationships between signals.
@@ IV-C4 Implementation of Moriano22
For detection, in each time window of length $\omega$, we compute the clustering similarity $s$ between the hierarchical clustering from training (computed once) and the hierarchical clustering computed during that specific time window. Here, similarity is used as a proxy of the distance between both hierarchical clusterings. We compute the similarity between hierarchical clustering using the CluSim method [61] based on the work by Gates et al. [62]. The similarity value provided by this method, $s$, lies in the range $[0,1]$, where 0 implies maximally dissimilar clusters, and 1 corresponds to identical clusters. Therefore, we report the anomaly score (or probability of detection) as $1-s$.
IV-C5 Algorithms’ Hyperparameter Setup
Matrix Correlation Distribution and Matrix Correlation Correlation algorithms do not require hyperparameters. They report anomaly scores based on $p$-values from the Mann-Whitney U test and Spearman correlation, respectively. For Ganesan17, we used the similar values as in Ganesan et al. [23], i.e., we set both $eps$ and $min\_samples$ to 1 to ensure a reasonable similarity between time series to be consider as part of the same cluster and allow individual time series to be part of its own cluster, respectively. For Moriano22, we used the sames values as in Moriano et al. [15], i.e., setting $r=-5.0$ and $\alpha=0.9$. Here, $r$ is a scaling parameter that defines the relative importance of memberships at different levels of the hierarchy. That is, the larger $r$, the more emphasis on comparing lower levels of the dendrogram (zoom in). In addition, $\alpha$ is a parameter that controls the influence of hierarchical clusterings with shared lineages. That is, the larger $\alpha$, the further the process will explore from the focus data element, so more of the cluster structure is taken into account into the comparison.
IV-D Dataset